# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ 8 Conclusion
The subscription business model, and its reliance on the "free trial with automatic renewal," has become a cornerstone of the modern digital economy. While the traditional economic rationale for trial offers centers on consumer learning, this paper has developed a complementary theoretical framework focused on a different, equally powerful friction: limited attention. We have demonstrated that when consumers are rationally inattentive and their ability to attend to future tasks decays over time, a rich set of strategic behaviors and market outcomes observed in subscription markets can be explained.
Our model is built upon a fundamental trade-off faced by the firm. On one side, extending a trial’s length erodes consumer attention, increasing the rate of accidental renewals and boosting "inattentive revenue." On the other side, sophisticated consumers anticipate this cognitive burden, which lowers their ex-ante utility and makes the offer less attractive, creating a countervailing "IR-Slack." We have shown that this tension naturally resolves to a finite, interior optimal trial length, providing a clear rationale for the lengthy but not infinite trial periods prevalent in the market.
The core predictions of our model, sharpened by standard and empirically-motivated assumptions, offer a new perspective on contract design and marketing strategy. We established a strategic complementarity between trial length and renewal price, predicting that firms using longer trials to leverage inattention will also set higher prices to maximize extraction from this channel. We extended the model to paid trials and articulated a "dilution principle," showing that upfront fees (which screen consumers) and long trials (which exploit inattention) act as strategic substitutes. Furthermore, our framework provides a powerful tool for policy analysis. We showed that "click-to-cancel" regulations are likely to shorten trial periods but will lead to higher prices if demand from loyal subscribers is inelastic.
@@
This paper represents a foundational step in exploring the role of attention in subscription contracts, and several promising avenues for future research remain open.
First, our analysis is conducted in a monopoly setting. Introducing competition would be a crucial and natural extension. How do firms design their contracts when consumers can choose between multiple subscription offers? Would competitive pressure force firms to offer shorter, more transparent trials, or could it perversely lead to a "race to the bottom," where firms compete on who can more effectively design contracts to exploit consumer inattention? The strategic interaction in such a setting would be complex and could yield valuable insights into market structure and consumer welfare.
@@
Third, the attention decay process itself could be modeled in greater detail. Our parameter $\beta$ is exogenous. One could imagine models where it is endogenous, for example, where consumers can invest in "attention capital" (e.g., by adopting organizational habits) to lower their personal decay rate. Moreover, firms could engage in dynamic activities during the trial period. For instance, sending engaging marketing emails might keep the product top-of-mind (effectively raising $\lambda$), while strategic silence might be used to deepen the attention decay.
Finally, the predictions of our model are empirically testable, and we believe this is a particularly fruitful path for future work. The predicted positive correlation between trial lengths and renewal prices, and the negative correlation between introductory prices and trial lengths, can be tested using market data from subscription services. Field experiments, conducted in partnership with firms, could be designed to randomly assign new users to different contract structures to causally identify the effects on sign-up, renewal, and long-term retention. Laboratory experiments could be used to validate the core behavioral mechanism of attention decay. Lastly, the model is suitable for structural estimation, which could use observational data on market contracts and consumer choices to recover the underlying parameters of the model, such as the distribution of consumer valuations and the attention parameters $\lambda_{0}$ and $\beta$. Such an estimation would allow for quantitative welfare analysis of various policy proposals.
In conclusion, by placing limited attention at the center of the analysis, this paper offers a new and, we believe, valuable perspective on the economics of the rapidly growing subscription industry, with important implications for both management practice and public policy.
@@
A.1 Proof of Lemma 1 (Optimal monitoring)
For any consumer whose valuation is below the renewal price, $v<P$, the cancellation reminder decision is
$\min_{q\in[0,1]}\;L(q;P,\lambda)\;=\;(1-q)P\;+\;\frac{1}{\lambda}\,H(q),\qquad H(q)=q\ln q+(1-q)\ln(1-q),$
with cost parameter $\lambda=\lambda(T)=\lambda_{0}/(1+\beta T)$.
First, we compute the first and second derivatives of the entropy term:
$H^{\prime}(q)=\ln\!\Bigl{(}\tfrac{q}{1-q}\Bigr{)},\qquad H^{\prime\prime}(q)=\frac{1}{q(1-q)}>0\quad\text{for }q\in(0,1).$
Hence $H(q)$ is strictly convex; scaling by the positive constant $1/\lambda$ preserves convexity. The linear loss $(1-q)P$ is also convex, so the entire objective $L(q;P,\lambda)$ is strictly convex on $(0,1)$.
Therefore any stationary point is the unique global minimizer.
Next, the derivative of $L$ is
$\frac{\partial L}{\partial q}=-P+\frac{1}{\lambda}\,H^{\prime}(q)=-P+\frac{1}{\lambda}\,\ln\!\Bigl{(}\tfrac{q}{1-q}\Bigr{)}.$
Setting $\partial L/\partial q=0$ yields
$\ln\!\Bigl{(}\tfrac{q}{1-q}\Bigr{)}=\lambda P\;\Longrightarrow\;\frac{q}{1-q}=e^{\lambda P}\;\Longrightarrow\;q^{*}(P,\lambda)=\frac{1}{1+e^{-\lambda P}}.$
Because $L$ is strictly convex and $q^{*}\in(0,1)$, expression (A.1 Proof of Lemma 1 (Optimal monitoring)) is the unique minimizer. Next, define $q^{*}\equiv q^{*}(P,\lambda)$. Using $q^{*}(1-q^{*})=e^{-\lambda P}/(1+e^{-\lambda P})^{2}$:
$\frac{\partial q^{*}}{\partial P}=\lambda q^{*}(1-q^{*})>0,\qquad\frac{\partial q^{*}}{\partial\lambda}=P\,q^{*}(1-q^{*})>0.$
Sensitivity to trial length $T$.
Since $\lambda(T)=\lambda_{0}/(1+\beta T)$ with $\beta>0$,
$\frac{\partial q^{*}}{\partial T}=\frac{\partial q^{*}}{\partial\lambda}\;\frac{d\lambda}{dT}=P\,q^{*}(1-q^{*})\;\Bigl{(}-\frac{\beta\lambda_{0}}{(1+\beta T)^{2}}\Bigr{)}<0.$
Boundary checks.
At $q=0$ the derivative of $L$ is $-P-\infty<0$; at $q=1$ it is $-P+\infty>0$. Hence the unique stationary point in $(0,1)$ is indeed a minimum, and no corner solution can dominate it.
Equations (A.1 Proof of Lemma 1 (Optimal monitoring))–(Sensitivity to trial length $T$.) establish
$q^{*}(P,\lambda)=\frac{1}{1+\exp(-\lambda P)},\quad\partial_{P}q^{*}>0,\;\partial_{\lambda}q^{*}>0,\;\partial_{T}q^{*}<0,$
thus completing the proof. ∎
@@ A.1 Proof of Lemma 1 (Optimal monitoring)
For a fixed renewal price $P>0$ recall that $q^{*}(P,\lambda)=\bigl{[}1+\exp(-\lambda P)\bigr{]}^{-1}$ from Lemma 1 and that $\lambda(T)=\lambda_{0}/(1+\beta T)$ with $\beta>0$ (MA-MA-3). By the chain rule
$\frac{\partial q^{*}}{\partial T}\;=\;\frac{\partial q^{*}}{\partial\lambda}\;\frac{d\lambda}{dT}.$
Step 1. Sign of $\partial q^{*}/\partial\lambda$.
Differentiate the closed-form expression:
[…]
because $q^{*}\in(0,1)$ for any $\lambda,P>0$.
Step 2. Sign of $d\lambda/dT$.
From $\lambda(T)=\lambda_{0}(1+\beta T)^{-1}$:
$\frac{d\lambda}{dT}=-\frac{\beta\lambda_{0}}{(1+\beta T)^{2}}\;<\;0\quad\text{for }T\geq 0,\;\beta>0.$
@@
Multiplying (LABEL:eq:A.2.2) and (LABEL:eq:A.2.3) inside (A.2. Proof of Proposition 1 (Attention vs. Trial length)) yields
$\boxed{\;\dfrac{\partial q^{*}(P,\lambda(T))}{\partial T}<0\;}\qquad\text{for all }P>0,\;T\geq 0,\;\beta>0.$
Strict inequality holds because both factors are strictly non-zero. $\blacksquare$
@@ A.3. Proof of Proposition 2 (Monotonicity of inattentive revenue)
Fix $P>0$. Definition (LABEL:eq:IR) gives
[…]
@@
From Proposition 1, $\partial q^{*}/\partial T<0$; hence $-\partial q^{*}/\partial T>0$.
* If $F(P)>0$ (i.e. there exists a positive mass of $v<P$), the factors $P$, $F(P)$, and $-\partial q^{*}/\partial T$ are all strictly positive, so
@@ A.4. Proof of Proposition 3 (Interior optimal $T^{*}$)
Throughout this proof $P>0$ is fixed and Assumptions MA-1, MA-3 ($\beta>0$) and C1(i) ($f(0)>0$) hold.
Derivative of profit with respect to $T$.
From (3),
@@
$\frac{\partial\Pi}{\partial T}(T,P)=P\,\frac{\partial\text{IR}}{\partial T}(T,P)\;-\;\text{IR-Slack}(T,P),$
exactly as stated in the proposition.
Sign near $T=0$.
Using (LABEL:eq:IR) and Lemma 1, $\partial q^{*}/\partial T=-\beta\lambda_{0}P\,q^{*}(1-q^{*})/(1+\beta T)^{2}.$ Hence, at $T=0$
[…]
Since $\lambda_{0}>0$ is finite, $q^{*}_{0}=1/(1+e^{-\lambda_{0}P})$ is strictly less than 1 for every $P>0$, so $q^{*}_{0}(1-q^{*}_{0})>0$. Therefore $\partial\text{IR}/\partial T|_{0}>0$.
At $T=0$ the IR-Slack term is
[…]
Because $q^{*}_{0}\to 1$ as $P\to 0$, we have $(-H(q^{*}_{0}))\sim(1-q^{*}_{0})\to 0$. Combining with (Sign near $T=0$.), we can choose a price interval $(0,\bar{P}]$ (any small $\bar{P}>0$ suffices) such that
[…]
Sign for large $T$.
As $T\to\infty$:
$\lambda(T)=\frac{\lambda_{0}}{1+\beta T}\longrightarrow 0,\quad q^{*}(P,\lambda(T))\longrightarrow\tfrac{1}{2},\quad-H\bigl{(}q^{*}\bigr{)}\longrightarrow\ln 2.$
[…]
Derivative of IR. Using $\partial q^{*}/\partial\lambda=P\,q^{*}(1-q^{*})\leq P/4$ and $|d\lambda/dT|=\beta\lambda_{0}/(1+\beta T)^{2}$:
[…]
From (Derivative of profit with respect to $T$.), for sufficiently large $T$ the positive Slack dominates the vanishing $P\,\partial\text{IR}/\partial T$; hence
$\exists\,\bar{T}<\infty\;:\;T>\bar{T}\;\Longrightarrow\;\frac{\partial\Pi}{\partial T}(T,P)<0.$
Existence of a root.
@@
The function $g(T):=\partial\Pi/\partial T$ is continuous in $T$ (derivatives of $q^{*}$ are continuous). By (LABEL:eq:A.4.4) $g(0)>0$; by (LABEL:eq:A.4.6) $g(T)<0$ for $T>\bar{T}$. Following Intermediate Value Theorem, there exists at least one $T^{*}\in(0,\bar{T})$ with $g(T^{*})=0$.
Step 4. Uniqueness.
Compute $dg/dT$:
$\frac{dg}{dT}=P\,\frac{\partial^{2}\text{IR}}{\partial T^{2}}-\frac{d\,\text{IR-Slack}}{dT}.$
* $P\,\partial^{2}\text{IR}/\partial T^{2}<0$ because $\partial q^{*}/\partial T$ is negative and decreasing in absolute value (Lemma 1 gives $\partial^{2}q^{*}/\partial T^{2}>0$). * $d\,\text{IR-Slack}/dT>0$ since $H(q^{*})$ is strictly increasing in $q^{*}$ and $q^{*}$ decreases in $T$.
Thus $dg/dT<0$ for all $T\geq 0$; $g(T)$ is strictly decreasing and can cross zero at most once. Therefore, the root found in Step 3 is unique.
Conclusion.
There exists a unique finite $T^{*}\in(0,\infty)$ satisfying $g(T^{*})=0$, i.e.
$P\,\frac{\partial\text{IR}}{\partial T}(T^{*},P)\;=\;\text{IR-Slack}(T^{*},P),$
which completes the proof of Proposition 3. ∎

### v3

@@ 8 Conclusion
The subscription business model, and its reliance on the "free trial with automatic renewal," has become a cornerstone of the modern digital economy. While the traditional economic rationale for trial offers centers on consumer learning, here we develop a complementary theoretical framework focused on a different, equally powerful friction: limited attention. We have demonstrated that when consumers are rationally inattentive and their ability to attend to future tasks decays over time, a rich set of strategic behaviors and market outcomes observed in subscription markets can be explained.
Our model is built upon a crucial trade-off faced by the firm. On the one hand, extending a trial’s length erodes consumer attention, increasing the rate of accidental renewals, and boosting "inattentive revenue". On the other hand, sophisticated consumers anticipate this cognitive burden, which lowers their ex-ante utility and makes the offer less attractive, creating a countervailing "IR-Slack." We have shown that this tension naturally resolves to a finite, interior optimal trial length, providing a clear rationale for the lengthy but not infinite trial periods prevalent in the market.
The core predictions of our model, logically follow from standard and empirically motivated assumptions, offer a new perspective on contract design and pricing strategy. We established a strategic complementarity between trial length and renewal price, predicting that firms that use longer trials to exploit inattention will also set higher prices to maximize extraction from this channel. We extended the model to paid trials and articulated a "dilution principle," showing that upfront fees (which screen consumers) and long trials (which exploit inattention) act as strategic substitutes. Furthermore, our framework provides a powerful tool for policy analysis. We showed that "click-to-cancel" regulations are likely to shorten trial periods but will lead to higher prices if demand from loyal subscribers is inelastic.
@@
Our paper represents a foundational step in exploring the role of attention in subscription contracts, and several promising avenues for future research remain open.
First, our analysis is conducted in a monopoly setting. Introducing competition would be a crucial and natural extension. How do firms design their contracts when consumers can choose between multiple subscription offers? Would competitive pressure force firms to offer shorter, more transparent trials, or could it perversely lead to a "race to the bottom," where firms compete on who can more effectively design contracts to exploit consumer inattention? The strategic interaction in such a setting would be complex and could yield valuable insights into market structure and consumer welfare. Our framework provides a useful tool to analyze this trade-off between competing on transparency versus competing on obfuscation.
@@
Third, the attention decay process itself could be modeled in greater detail. Our parameter $\beta$ is exogenous. One could imagine models where it is endogenous, for example, where consumers can invest in "attention capital" (e.g., by adopting organizational habits) to lower their personal decay rate. Moreover, firms could engage in dynamic activities during the trial period. For instance, sending engaging reminder emails might keep the product top-of-mind (effectively raising $\tau$), while strategic silence might be used to deepen the attention decay.
Finally, the predictions of our model are empirically testable and we believe that this is a particularly fruitful path for future work. The predicted positive correlation between trial lengths and renewal prices, and the negative correlation between introductory prices and trial lengths, can be tested using market data from subscription services. Field experiments, conducted in partnership with firms, could be designed to randomly assign new users to different contract structures to causally identify the effects on sign-up, renewal, and long-term retention. Recent regulatory actions also create interesting natural experiments, where the geographic variations in trial and auto-renewal regulations can be leveraged for identification. Lastly, the model is suitable for structural estimation, which could use observational data on market contracts and consumer choices to recover the underlying parameters of the model, such as the distribution of consumer valuations and the attention parameters $\tau_{0}$ and $\beta$. Such an estimation would allow for quantitative welfare analysis of various policy proposals.
In conclusion, by placing limited attention at the center of the analysis, our framework as presented here offers a new and, we believe, valuable perspective on the economics of the rapidly growing subscription industry, with important implications for both management practice and public policy.
@@
A.1. Proof of Lemma 1 (Optimal Monitoring)
A consumer with valuation $v<P$ chooses a cancellation probability $q\in[0,1]$ to solve the cost-minimization problem laid out in Equation (3.5):
$\min_{q\in[0,1]}\quad\underbrace{\bigl(1-q\bigr)P}_{\text{Expected Monetary Loss}}\;+\;\underbrace{\frac{1}{\tau(T)}\,\mathcal{I}(q\parallel 1/2)}_{\text{Cognitive Cost}},$
where $\mathcal{I}(q\parallel 1/2)=[q\ln q+(1-q)\ln(1-q)]+\ln 2$ is the mutual information.
Step 1: Simplification of the Objective.
As the term $\ln 2/\tau(T)$ is constant with respect to the choice variable $q$, it does not influence the optimal solution. The consumer’s problem is therefore equivalent to solving:
$\min_{q\in[0,1]}\quad L(q)=(1-q)P+\frac{1}{\tau(T)}\Bigl[q\ln q+(1-q)\ln(1-q)\Bigr].$
Step 2: Convexity and Existence of a Unique Minimum.
The objective function $L(q)$ is strictly convex in $q$ on the domain $(0,1)$ because its second derivative is strictly positive:
$\frac{\partial^{2}L}{\partial q^{2}}=\frac{1}{\tau}\frac{d^{2}}{dq^{2}}\left[q\ln q+(1-q)\ln(1-q)\right]=\frac{1}{\tau q(1-q)}>0.$
Strict convexity guarantees that any stationary point is a unique global minimum.
Step 3: First-Order Condition.
The first-order condition for an interior solution is $\partial L/\partial q=0$:
$P+\frac{1}{\tau}\ln\left(\frac{q}{1-q}\right)=0.$
Solving for $q$ yields $\ln(q/(1-q))=\tau P$, which gives the stated logistic form for the optimal cancellation probability.
Step 4: Boundary behavior.
As $q\to 0^{+}$ the derivative tends to $P-\infty<0$; as $q\to 1^{-}$ it tends to $P+\infty>0$. Thus the interior stationary point derived above is indeed the global minimum.
Step 5: Comparative statics.
Using the closed form and writing $q^{*}\equiv q^{*}(P,\tau)$:
$\frac{\partial q^{*}}{\partial P}$ $=\tau q^{*}(1-q^{*})\,>0,$
$\frac{\partial q^{*}}{\partial\tau}$ $=P\,q^{*}(1-q^{*})\,>0.$
With $\tau(T)=\tau_{0}/(1+\beta T)$ and $\beta>0$,
$\frac{\partial q^{*}}{\partial T}=\frac{\partial q^{*}}{\partial\tau}\frac{d\tau}{dT}=-\,\frac{\beta\tau_{0}P}{(1+\beta T)^{2}}\,q^{*}(1-q^{*})<0.$
@@ A.1. Proof of Lemma 1 (Optimal Monitoring)
For a fixed renewal price $P>0$ recall that $q^{*}(P,\tau)=\bigl[1+\exp(-\tau P)\bigr]^{-1}$ from Lemma 1 and that $\tau(T)=\tau_{0}/(1+\beta T)$ with $\beta>0$ (MA‑MA-3). By the chain rule
$\frac{\partial q^{*}}{\partial T}\;=\;\frac{\partial q^{*}}{\partial\tau}\;\frac{d\tau}{dT}.$
Step 1. Sign of $\partial q^{*}/\partial\tau$.
Differentiate the closed‑form expression:
$\frac{\partial q^{*}}{\partial\tau}=\frac{\partial}{\partial\tau}\Bigl[1+\exp(-\tau P)\Bigr]^{-1}=P\,q^{*}(1-q^{*})\;>\;0,$
because $q^{*}\in(0,1)$ for any $\tau,P>0$.
Step 2. Sign of $d\tau/dT$.
From $\tau(T)=\tau_{0}(1+\beta T)^{-1}$:
$\frac{d\tau}{dT}=-\frac{\beta\tau_{0}}{(1+\beta T)^{2}}\;<\;0\quad\text{for }T\geq 0,\;\beta>0.$
@@
$\boxed{\;\dfrac{\partial q^{*}(P,\tau(T))}{\partial T}<0\;}\qquad\text{for all }P>0,\;T\geq 0,\;\beta>0.$
Strict inequality holds because both factors are strictly non‑zero. ∎
@@ A.3. Proof of Proposition 2 (Monotonicity of inattentive revenue)
Fix $P>0$. Definition (2) gives
$\text{IR}(T,P)=P\int_{0}^{P}\bigl[1-q^{*}(P,\tau(T))\bigr]\,f(v)\,dv=P\,F(P)\,\bigl[1-q^{*}(P,\tau(T))\bigr],$
@@
From Proposition 1, $\partial q^{*}/\partial T<0$; hence $\partial q^{*}/\partial T>0$.
* If $F(P)>0$ (i.e. there exists a positive mass of $v<P$), the factors $P$, $F(P)$, and $\partial q^{*}/\partial T$ are all strictly positive, so
@@ A.4. Proof of Proposition 3 (Interior optimal $T^{*}$)
Throughout this proof, $P>0$ is fixed and we assume MA-1, MA-3 ($\beta>0$), and A2 (positive mass at low valuations).
Step 1: Set up the Lagrangian.
The firm’s problem with a binding participation constraint is to maximize the Lagrangian $\mathcal{L}=\Pi(T,P)-\mu U(T,P)$, where $\mu>0$. The first-order condition with respect to $T$ is $\frac{\partial\mathcal{L}}{\partial T}=0$.
Step 2: Characterize the FOC.
@@
The FOC can be written as $\frac{\partial\text{IR}}{\partial T}=\mu\cdot\text{IR-Slack}(T,P)$. Let $g(T):=\frac{\partial\text{IR}}{\partial T}(T,P)-\mu\cdot\text{IR-Slack}(T,P)$. The optimal $T^{*}$ must satisfy $g(T^{*})=0$.
Step 3: Behavior near $T=0$.
At $T=0$, both $\frac{\partial\text{IR}}{\partial T}|_{T=0}$ and $\text{IR-Slack}(0,P)$ are positive and finite. For an interior solution to exist, the firm must have an incentive to increase the trial length from zero, which requires $g(0)>0$. This condition holds if the marginal revenue from increased inattention initially outweighs the marginal harm to consumer utility.
Step 4: Behavior as $T\to\infty$.
We prove that for a sufficiently large $T$, $g(T)<0$ by analyzing the asymptotic behavior of its components as $T\to\infty$ (which corresponds to $\tau\to 0$). We use second-order Taylor expansions around $\tau=0$. The optimal cancellation probability is $q^{*}=(1+\exp(-\tau P))^{-1}=\frac{1}{2}+\frac{\tau P}{4}+O(\tau^{3})$.
First, we analyze the marginal benefit term, $\frac{\partial\text{IR}}{\partial T}=PF(P)(-\frac{\partial q^{*}}{\partial\tau}\frac{d\tau}{dT})$.
The components are: $\frac{d\tau}{dT}=-\frac{\beta\tau_{0}}{(1+\beta T)^{2}}=-\frac{\beta}{\tau_{0}}\tau^{2}$. $\frac{\partial q^{*}}{\partial\tau}=Pq^{*}(1-q^{*})=P\left(\frac{1}{4}-\frac{(\tau P)^{2}}{16}+O(\tau^{4})\right)=\frac{P}{4}+O(\tau^{2})$.
Thus, the marginal benefit term has the asymptotic form:
$\frac{\partial\text{IR}}{\partial T}=PF(P)\left(-\frac{P}{4}+O(\tau^{2})\right)\left(-\frac{\beta}{\tau_{0}}\tau^{2}\right)=\frac{\beta P^{2}F(P)}{4\tau_{0}}\tau^{2}+O(\tau^{4}).$
Next, we analyze the marginal cost term, $\mu\cdot\text{IR-Slack}(T,P)=\mu\frac{\beta}{\tau_{0}}F(P)\,\mathcal{I}(q^{*}\parallel 1/2)$. The mutual information $\mathcal{I}(q\parallel 1/2)$ has a Taylor expansion around $q=1/2$ of $\mathcal{I}(q)=2(q-1/2)^{2}+O((q-1/2)^{4})$. Substituting $q^{*}-1/2=\tau P/4+O(\tau^{3})$:
$\mathcal{I}(q^{*}\parallel 1/2)=2\left(\frac{\tau P}{4}\right)^{2}+O(\tau^{4})=\frac{\tau^{2}P^{2}}{8}+O(\tau^{4}).$
Thus, the IR-Slack has the asymptotic form:
$\text{IR-Slack}(T,P)=\frac{\beta F(P)}{\tau_{0}}\left(\frac{\tau^{2}P^{2}}{8}+O(\tau^{4})\right).$
Combining these terms, the asymptotic form of $g(T)$ for small $\tau$ is:
$g(T)=\frac{\beta P^{2}F(P)}{4\tau_{0}}\tau^{2}-\mu\frac{\beta P^{2}F(P)}{8\tau_{0}}\tau^{2}+O(\tau^{4})=\frac{\beta P^{2}F(P)\tau^{2}}{8\tau_{0}}(2-\mu)+O(\tau^{4}).$
At an interior optimum, the firm must be extracting value, which implies the shadow value of consumer utility, $\mu$, must be sufficiently large. Under standard conditions ensuring a non-trivial trade-off, $\mu>2$. Therefore, the term $(2-\mu)$ is negative, which implies $g(T)<0$ for sufficiently large $T$.
Step 5: Existence by Intermediate Value Theorem.
@@
Since $g(T)$ is a continuous function, $g(0)>0$, and $g(T)<0$ for large $T$, the Intermediate Value Theorem guarantees that there exists at least one $T^{*}\in(0,\infty)$ such that $g(T^{*})=0$.
Step 6: Uniqueness.
Uniqueness of the optimum is guaranteed if the firm’s objective, $\mathcal{L}(T)$, is strictly concave in $T$. We examine its second derivative:
$\frac{\partial^{2}\mathcal{L}}{\partial T^{2}}=\frac{\partial^{2}\Pi}{\partial T^{2}}-\mu\frac{\partial^{2}U}{\partial T^{2}}.$
Let’s analyze the components:
  • $\frac{\partial^{2}\Pi}{\partial T^{2}}=\frac{\partial^{2}\text{IR}}{\partial T^{2}}$. This term represents the change in the marginal benefit of inattentive revenue. Standard assumptions of diminishing returns ensure this is negative.
  • $\frac{\partial^{2}U}{\partial T^{2}}=\frac{\partial}{\partial T}\left(-\text{IR-Slack}(T,P)\right)=-\frac{d(\text{IR-Slack})}{dT}$. As $T$ increases, $\tau$ decreases, and $q^{*}$ moves closer to $1/2$. The mutual information $\mathcal{I}(q^{*}\parallel 1/2)$ is a convex function of $q^{*}$ with a minimum at $q^{*}=1/2$. Since $q^{*}$ is a decreasing function of $T$, $\mathcal{I}(q^{*}\parallel 1/2)$ is also a decreasing function of $T$. This implies $\frac{d(\text{IR-Slack})}{dT}<0$, and therefore $\frac{\partial^{2}U}{\partial T^{2}}>0$.
Substituting these back into the second derivative of the Lagrangian:
$\frac{\partial^{2}\mathcal{L}}{\partial T^{2}}=\underbrace{\frac{\partial^{2}\Pi}{\partial T^{2}}}_{\text{Negative}}-\underbrace{\mu}_{>0}\underbrace{\frac{\partial^{2}U}{\partial T^{2}}}_{\text{Positive}}<0.$
Since the Lagrangian is strictly concave in $T$, there can be at most one value of $T$ that satisfies the first-order condition. This guarantees the uniqueness of the optimal trial length $T^{*}$.